Mike drew a painting on a rectangular paper size 59 by 43 centimeter. Later he places the painting in a rectangular frame of which a uniform space is left around the painting. If the perimeter of the base of the frame is 256 centimeters. what is the width of the space that surrounds the painting?
Answer by ankor@dixie-net.com(12701)

Mike drew a painting on a rectangular paper size 59 by 43 centimeter. Later he places the painting in a rectangular frame of which a uniform space is left around the painting. If the perimeter of the base of the frame is 256 centimeters. what is the width of the space that surrounds the painting?
:
Let x = the width of the space around the 59 by 43 cm painting
:
The dimensions of the frame will be:
(59+2x) by (43+2x)
:
The perimeter of the frame will be:
2(59+2x) + 2(43+2x) = 256
:
Simplify equation divide by 2:
59 + 2x + 43 + 2x = 128
:
4x + 102 = 128
:
4x = 128 - 102
:
x = 26/4
:
x = 6.5" is the width of the area around the painting
:
:
Check solution using the perimeter:
Frame: (59+13) by (43+13)
2(72) + 2(56) =
144 + 112 = 256